OpenStudy (anonymous):

please help!! Francine and Cheryl received equal scores on a test made up of multiple choice questions and an essay. Francine got 34 multiple choice questions correct and received 14 points for her essay. Cheryl got 30 multiple choice questions correct and received 22 points for her essay. How many points was each multiple choice question worth?

OpenStudy (cwrw238):

let x be the number of points for a multiple choice question then FRancine got 34x + 14 points Cheryl got 30x + 22 points these are equal, so 34x + 14 = 30x + 22 solve for x

OpenStudy (anonymous):

soo add 30x+22?

OpenStudy (cwrw238):

34x + 14 = 30x + 22 -14 -14 34x = 30x + 8 -30x -30x 4x = 8 x = 8/4 = 2